# Fisher-Rao Gradient Flows of Linear Programs and State-Action Natural Policy Gradients

### v1

@@ 3.2. Estimating the regularization error
See 3.3
Proof.
Recall that by the central path property the Fisher-Rao gradient flow $(\mu_{t})_{t\geq 0}$ satisfies
$\mu_{t}=\operatorname*{arg\,max}\left\{c^{\top}\mu-t^{-1}D_{\operatorname{KL}}(\mu,\mu_{0}):\mu\in P\right\}.$
If we choose $\mu_{0}$ as the information projection according to (3.28) we obtain by the Pythagorean theorem that
@@
This shows that
$\mu_{t}=\operatorname*{arg\,max}\left\{c^{\top}\mu+t^{-1}H(\mu):\mu\in P\right\},$
i.e., that $\mu_{t}$ is the solution of the entropy regularized linear program (3.6). Now the claim follows from Theorem 3.2 and $D_{\operatorname{KL}}(\mu^{\star},\mu_{0})=H(\mu_{0})-H(\mu^{\star})\leq R_{H}$. ∎
@@ 3.3. Non-unique maximizers
Both Theorem 3.2 and Corollary 3.3 are formulated under the assumption that the linear program (3.1) admits a unique solution. This is satisfied for almost all cost vectors $c\in\mathbb{R}^{\mathbb{X}}$, however, it can be generalized to all costs. To this end, we use the following generalization of Lemma 3.8 where we defer the proof to the appendix as it follows similar arguments.
Lemma 3.11.
and further
To proceed like in the proof with a unique maximizer, we need to identify the limit of $\mu_{t}$ in $F^{\star}$. For linear objectives function the limit $\mu^{\star}$ is the information projection of $\mu_{0}$ to $F^{\star}$, see [5, Corollary 4.8]. We include a proof here for the sake of completeness.
@@
In words, the Fisher-Rao gradient flow converges to the information projection of $\mu_{0}$ to $F^{\star}$, i.e., it selects the optimizer which has the minimum KL-divergence from $\mu_{0}$.
@@
Consider Setting 3.1, assume that the linear program is non trivial, i.e., that $F^{\star}\neq P$, where $F^{\star}$ denotes the face of optimizers, and denote the information projection of $\mu_{0}$ to $F^{\star}$ by $\mu^{\star}\in F^{\star}$ and set
@@ 4.1. Compatible function approximation
In this subsection, and more precisely in Proposition 4.4, we describe the natural gradient direction as the minimizer of a linear least squares regression problem with features $\phi_{\theta}(x)=\nabla_{\theta}\log\mu_{\theta}(x)$, which is known as compatible function approximation in reinforcement learning. This can be used to approximate the natural gradient using samples from the distribution $\mu_{\theta}$.
@@
Further, by the chain rule, we have
$\partial_{t}\mu_{t}(x)=\partial_{t}\theta_{t}^{\top}\nabla_{\theta}\mu_{\theta_{t}}(x)=v_{t}^{\top}\nabla_{\theta}\mu_{\theta_{t}}(x).$
Using $\nabla^{\operatorname*{\operatorname{FR}}}f(\mu)=\nabla f(\mu)\odot\mu$ we conclude
@@
$=\mathbb{E}_{\mu_{\theta}}\left[\left(v_{t}^{\top}\nabla_{\theta}\log\mu_{\theta}(x)-\nabla f(\mu_{\theta})(x)\right)^{2}\right]$
$=L(v_{t},\theta).$
@@
is often called the approximation error. Examining the objective $L(w,\theta)$ and using Lemma 4.3 we see that the natural gradient flow minimizes the discrepancy between $\partial_{t}\mu_{t}$ and the Fisher-Rao gradient $\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})$. In this case, the evolution $\partial_{t}\mu_{t}$ is given by the orthogonal projection of the Fisher-Rao gradient onto the tangent space of the parametrized model. A similar property holds for any natural gradient defined using a Riemannian metric on the polytope [7, 52, 40].
@@
$\lVert\partial_{t}\mu_{t}-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}$ $=\min_{w\in\mathbb{R}^{p}}\left\lVert w^{\top}\nabla_{\theta}d_{\theta}-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\right\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}$
$=\min_{\nu\in T_{\theta}P}\left\lVert\nu-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\right\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}.$
@@ 4.2. Convergence of natural gradient flows
$\partial_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t})$ $=-\partial_{t}\phi(\mu_{t})-\partial_{t}\langle\nabla\phi(\mu_{t}),\mu^{\star}-\mu_{t}\rangle$
$=\langle\nabla^{2}\phi(\mu_{t})\partial_{t}\mu_{t},\mu_{t}-\mu^{\star}\rangle$
@@
$\leq\nabla f(\mu_{t})^{\top}(\mu_{t}-\mu^{\star})+\varepsilon_{t}\delta_{t}$
$\leq f(\mu_{t})-f(\mu^{\star})+\varepsilon_{t}\delta_{t},$
@@
In reinforcement learning, baselines are often used when estimating the natural policy gradient from samples to reduce the variance of the estimates [54]. This amounts to projecting the gradient of the objective to the tangent space of the model. In our setting, this corresponds to projecting $\nabla f(\mu)\odot\mu$ to the tangent space $TP$ with respect to the Fisher-Rao metric $g^{\operatorname*{\operatorname{FR}}}_{\mu}$, see also [51, Subsection 4.1.1]. In the special case $P=\Delta_{\mathbb{X}}$ the Fisher-Rao projection of $\nabla f(\mu)\odot\mu$ is given by $\nabla f(\mu)\odot\mu-\kappa\mu$, where $\kappa=\sum_{x}\nabla f(\mu)(x)$ and the corresponding compatible function approximation objective is given by
@@ 4.3. Global convergence for multi-player games
For a rich enough para-metrization Corollary 4.5 ensures that $(\mu_{\theta_{t}})_{t\geq 0}$ follows the Fisher-Rao gradient flow in which case Theorem 3.2 implies the linear convergence of the natural gradient flow (4.1). A common example is the softmax parametrization $\mu_{\theta}(x)\propto e^{\theta(x)}$. For multi-player games with suitable payoff structure, the dynamics of the individual players decouple [13], which allows us to provide a global convergence result for parametric models with exponentially fewer parameters than the softmax parametrization.
@@
Consider a differentiable parametrization of conditional probabilities $\{m_{\theta}:\theta\in\mathbb{R}^{p}\}=\operatorname{int}(\Delta_{\mathbb{X}}^{n})$, where $n\in\mathbb{N}$ and $\mathbb{X}$ is a finite set, and suppose that
$\operatorname{span}\left\{\partial_{\theta_{i}}m_{\theta}:i=1,\dots,p\right\}=T\Delta_{\mathbb{X}}^{n}\quad\text{for all }\theta\in\mathbb{R}^{p}.$ (4.10)
Define a corresponding parametric independence model as
@@ 5. Convergence of State-Action Natural Policy Gradients
They determine the reward as
$R(\pi)=\sum_{s\in\mathbb{S},a\in\mathbb{A}}r(s,a)d^{\pi}(s,a)=r^{\top}d^{\pi}.$ (5.5)
The set of state-action distributions has been characterized as a polytope in the seminal works of Cyrus Derman, see [19].
@@
where we write $R(\theta)=R(\pi_{\theta})$. Here $G(\theta)$ denotes a Gramian matrix with entries $G(\theta)_{ij}=g_{d_{\theta}}(\partial_{\theta_{i}}d_{\theta},\partial_{\theta_{j}}d_{\theta})$, where we write $d_{\theta}=d^{\pi_{\theta}}$ and $g_{d}$ denotes a Riemannian metric on the state-action polytope $\mathscrsfs{D}$. In this context, the matrix $G(\theta)$ is referred to as a preconditioner. Various choices have been suggested for $G(\theta)$, such as for example
@@
which is a weighted sum of Fisher-information matrices over the individual states [24, 9, 43] which has been studied extensively in the literature. We focus on the so-called state-action natural gradient given by the Fisher information matrix of the state-action distribution [35],
@@
This choice was observed to reduce the severity of plateaus and is closely connected to the trust region method known as relative entropy policy search (REPS) [42].
5.1. Convergence of state-action natural policy gradients
@@
Consider a finite discounted Markov decision process, suppose 5.2 and 5.3 hold, and consider a solution of the natural policy gradient flow (5.10) for $G=G_{\operatorname{M}}$. Then it holds that
@@
  • Expressive exponential families: For a feature map $\phi\colon\mathbb{S}\times\mathbb{A}\to\mathbb{R}^{p}$ and $\theta\in\mathbb{R}^{p}$ we consider the log-linear policy
    $\pi_{\theta}(a|s)\coloneqq\frac{e^{\theta^{\top}\phi(s,a)}}{\sum_{a^{\prime}}e^{\theta^{\top}\phi(s,a^{\prime})}}.$
    If $\operatorname{rank}\{\phi(s,a):s\in\mathbb{S},a\in\mathbb{A}\}=\lvert\mathbb{S}\rvert\cdot\lvert\mathbb{A}\rvert$ this provides a regular tabular parametrization, see [47, Remark 2.4]. In particular, this includes tabular softmax policies, where $\pi_{\theta}(a|s)\propto e^{\theta_{s,a}}$ which is the arguably most commonly studied policy class in policy gradient methods.
  • Escort transform: To reduce the plateaus of vanilla policy gradients when working with softmax policies the so-called escort transform given by
    $\pi_{\theta}(a|s)\coloneqq\frac{\lvert\theta_{s,a}\rvert^{p}}{\sum_{a^{\prime}}\lvert\theta_{s,a^{\prime}}\rvert^{p}},$
    for $p\geq 1$ was introduced in [32].
@@
Consider a finite discounted Markov decision process, suppose 5.2 and 5.3 hold, and consider a solution of the natural gradient flow (5.10) for a regular tabular parametrization and write $d_{t}=d_{\theta_{t}}$. Further, assume that there is a unique optimal policy $\pi^{\star}$ with state-action distribution $d^{\star}$ and optimal reward $R^{\star}$. Then we have
$D_{\operatorname{KL}}(d^{\star},d_{t})\leq D_{\operatorname{KL}}(d^{\star},d_{0})\exp\left(-\Delta(t-t_{0})+2t_{0}\Delta\log\left(\frac{t+t_{0}}{2t_{0}}\right)\right)$ (5.16)
and
$R^{\star}-R(\theta_{t})\leq\Delta D_{\operatorname{KL}}(d^{\star},d_{0})\exp\left(-\Delta(t-t_{0})+2t_{0}\Delta\log\left(\frac{t+t_{0}}{2t_{0}}\right)\right)$ (5.17)
for all $t\geq t_{0}$, where
$\begin{split}\Delta&\coloneqq\min\left\{\frac{R^{\star}-R(\pi)}{\lVert d^{\star}-d^{\pi}\rVert_{\operatorname*{\operatorname{TV}}}}:\begin{array}[]{c}\pi\text{ is deterministic and}\\ \text{agrees with }\pi^{\star}\text{ on all but one state}\end{array}\right\}>0,\end{split}$ (5.18)
and $N(d^{\star})$ denotes the neighboring vertices of $d^{\star}$ in the state-action polytope $\mathscrsfs{D}$ and
$\begin{split}t_{0}&\coloneqq\frac{2D_{\operatorname{KL}}(d^{\star},d_{0})}{\Delta\cdot\min\{d^{\star}(s):s\in\mathbb{S}\}}.\end{split}$ (5.19)
@@
Note that we can apply Theorem 3.2 and hence it remains to provide the explicit expressions for $\Delta$ and $t_{0}$. Note that since $\pi^{\star}$ is deterministic, we have $d^{\star}(s,a)=d^{\star}(s)$ if $a$ is optimal and $d^{\star}(s,a)=0$ otherwise. Hence, we have
$\min\{d^{\star}(s,a):s\in\mathbb{S},a\in\mathbb{A}\}=\min\{d^{\star}(s):s\in\mathbb{S}\}$
showing (5.19). For the explicit expression of $\Delta$, note that the neighboring vertices in $\mathscrsfs{D}$ and $\Delta_{\mathbb{A}}^{\mathbb{S}}$ correspond to each other [38]. Hence, $d^{\pi}$ is a neighbor of $d^{\star}$ if $\pi$ is deterministic and agrees with $\pi^{\star}$ on all but one state. ∎
@@
Much like the state-action natural policy gradient, Kakade’s natural policy gradient with exact gradient evaluations has been shown to convergence linearly without the need of entropy regularized setting [27]. Here, the discrete-time setting is studied and NPG is interpreted as soft policy iteration. This is used to show a convergence rate of $R^{\star}-R(\pi_{k})=O(e^{-ck})$ for any $c\in(0,\Delta_{\operatorname{K}})$, where
$\Delta_{\operatorname{K}}\coloneqq-(1-\gamma)^{-1}\max\left\{A^{\star}(s,a):a\neq a_{s}^{\star}\right\}\geq\Delta.$
Indeed, by the performance difference lemma, we have
@@
Note that since $d\in N(d^{\star})$ and since $A^{\star}(s,a_{s}^{\star})=0$ we have $\pi(a_{0}|s_{0})=1$ with $a_{0}\neq a_{s_{0}}^{\star}$ for and $\pi(a_{s}^{\star}|s)=1$ for $s\neq s_{0}$. Now we have $d^{T}A^{\star}=d(s_{0})A^{\star}(s_{0},a_{0})\leq 0$. Further, we estimate
@@
$\Delta\leq-(1-\gamma)\max\left\{A^{\star}(s,a):a\neq a_{s}^{\star}\right\}=\Delta_{\operatorname{K}}.$
Hence, the best-known convergence rate of Kakade’s NPG is faster compared to the rate we provide for the state-action natural policy gradient. However, in the general case, no lower bounds are known for either gradient methods. In our computational example, both converge at the same exponential rate $O(e^{-\Delta_{\operatorname{K}}t})$ even if $\Delta<\Delta_{\operatorname{K}}$.
5.2. A computational example
We use the following example proposed by Kakade [24] and which was also used in [9, 35]. Computer code for all experiments is made available under https://github.com/muellerjohannes/fisher-rao-GFs-LPs. We consider an MDP with two states $s_{1},s_{2}$ and two actions $a_{1},a_{2}$, with the transitions and instantaneous rewards shown in Figure 2.

### v2

@@ 3.2. Estimating the regularization error
Proof of Corollary 3.3.
By the central path property the Fisher-Rao gradient flow $(\mu_{t})_{t\geq 0}$ satisfies $\mu_{t}=\operatorname*{arg\,max}\left\{c^{\top}\mu-t^{-1}D_{\operatorname{KL}}(\mu,\mu_{0}):\mu\in P\right\}$. If we choose $\mu_{0}$ as the information projection according to (3.24) the Pythagorean theorem yields
@@
This shows that $\mu_{t}=\operatorname*{arg\,max}\left\{c^{\top}\mu+t^{-1}H(\mu):\mu\in P\right\}$, i.e., that $\mu_{t}$ is the solution of the entropy regularized linear program (3.6). Now the claim follows from Theorem 3.2 and $D_{\operatorname{KL}}(\mu^{\star},\mu_{0})=H(\mu_{0})-H(\mu^{\star})\leq R_{H}$. ∎
@@ 3.3. Non-unique maximizers
Both Theorem 3.2 and Corollary 3.3 are formulated under the assumption that the linear program (3.1) admits a unique solution. This is satisfied for almost all costs $c\in\mathbb{R}^{\mathbb{X}}$, however, it can be generalized to all costs.
To proceed like in the proof with a unique maximizer, we need to identify the limit of $\mu_{t}$ in $F^{\star}$. For linear objective functions the limit $\mu^{\star}$ is the information projection of $\mu_{0}$ to $F^{\star}$, see [5, Corollary 4.8]. We include a proof here for the sake of completeness.
@@
In words, the Fisher-Rao gradient flow converges to the information projection of $\mu_{0}$ to $F^{\star}$, i.e., it selects the optimizer that has the minimum KL-divergence from $\mu_{0}$.
@@
Consider Setting 3.1, assume that the linear program is non-trivial, i.e., that $F^{\star}\neq P$, where $F^{\star}$ denotes the face of optimizers, and denote the information projection of $\mu_{0}$ to $F^{\star}$ by $\mu^{\star}\in F^{\star}$ and set
@@ 4.1. Compatible function approximation
In this subsection, and more precisely in Proposition 4.4, we describe the natural gradient direction as the minimizer of a linear least squares regression problem with features $\phi_{\theta}(x)=\nabla_{\theta}\log\mu_{\theta}(x)$. This can be used to estimate the natural gradient from samples drawn from $\mu_{\theta}$.
In the context of reinforcement learning similar techniques, albeit for a different notion of natural gradient, have been developed under the name compatible function approximation [55, 26, 2].
@@
Further, by the chain rule, we have $\partial_{t}\mu_{t}=\partial_{t}\theta_{t}^{\top}\nabla_{\theta}\mu_{\theta_{t}}(x)=v_{t}^{\top}\nabla_{\theta}\mu_{\theta_{t}}(x)$. Using $\nabla^{\operatorname*{\operatorname{FR}}}f(\mu)=\nabla f(\mu)\odot\mu$ we conclude
@@
$=\mathbb{E}_{\mu_{\theta}}\left[\left(v_{t}^{\top}\nabla_{\theta}\log\mu_{\theta}(x)-\nabla f(\mu_{\theta})(x)\right)^{2}\right]=L(v_{t},\theta).$
The distance between $\partial_{t}\mu_{t}$ and the Fisher-Rao gradient $\nabla^{\textup{FR}}_{P}f(\mu_{t})$ is up to a remainder term given by the least squares loss $L(v_{t},\theta_{t})$, where $v_{t}=\partial_{t}\theta_{t}$. The natural gradient is designed such that $\partial_{t}\mu_{t}$ is close to $\nabla^{\textup{FR}}_{P}f(\mu_{t})$ [6] and hence we can minimize the least squares loss $L(v,\theta_{t})$ with respect to $v$ in order to approximate the natural gradient $v_{t}\approx F(\theta_{t})^{+}\nabla f(\theta_{t})$. An important benefit of this formulation is that it can be used to estimate the natural gradient from data distributed according to $\mu_{\theta_{t}}$. We make this relation between the minimization of $L$ and the natural gradient explicit.
@@
is can be interpreted as an approximation error. Note, however, that the precise nature of the least square loss $L$ is different from the one well-known in reinforcement learning as we discuss in more detail in Remark 5.6. Examining the objective $L(w,\theta)$ and using Lemma 4.3 we see that the natural gradient flow minimizes the discrepancy between $\partial_{t}\mu_{t}$ and the Fisher-Rao gradient $\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})$. In this case, the evolution $\partial_{t}\mu_{t}$ is given by the orthogonal projection of the Fisher-Rao gradient onto the tangent space of the parametrized model. A similar property holds for any natural gradient defined using a Riemannian metric on the polytope [7, 57, 43].
@@
$\lVert\partial_{t}\mu_{t}-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}$ $=\min_{w\in\mathbb{R}^{p}}\left\lVert w^{\top}\nabla_{\theta}d_{\theta}-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\right\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}\min_{\nu\in T_{\theta}P}\left\lVert\nu-\nabla^{\operatorname*{\operatorname{FR}}}_{P}f(\mu_{t})\right\rVert_{g^{\operatorname*{\operatorname{FR}}}_{\mu_{t}}}.$
@@ 4.2. Convergence of natural gradient flows
$\partial_{t}D_{\operatorname{KL}}(\mu^{\star},\mu_{t})$ $=-\partial_{t}\phi(\mu_{t})-\partial_{t}\langle\nabla\phi(\mu_{t}),\mu^{\star}-\mu_{t}\rangle=\langle\nabla^{2}\phi(\mu_{t})\partial_{t}\mu_{t},\mu_{t}-\mu^{\star}\rangle$
@@
$\leq\nabla f(\mu_{t})^{\top}(\mu_{t}-\mu^{\star})+\varepsilon_{t}\delta_{t}\leq f(\mu_{t})-f(\mu^{\star})+\varepsilon_{t}\delta_{t},$
@@
In reinforcement learning, baselines are often used when estimating the natural policy gradient from samples to reduce the variance of the estimates [59]. This amounts to projecting the gradient of the objective to the tangent space of the model. In our setting, this corresponds to projecting $\nabla f(\mu)\odot\mu$ to the tangent space $TP$ with respect to the Fisher-Rao metric $g^{\operatorname*{\operatorname{FR}}}_{\mu}$, see also [54, Subsection 4.1.1]. In the special case $P=\Delta_{\mathbb{X}}$ the Fisher-Rao projection of $\nabla f(\mu)\odot\mu$ is given by $\nabla f(\mu)\odot\mu-\kappa\mu$, where $\kappa=\sum_{x}\nabla f(\mu)(x)$ and the corresponding compatible function approximation objective is given by
@@ 4.3. Global convergence for multi-player games
With function approximation Corollary 4.7 ensures sublinear convergence $O(\frac{1}{t})$ up to a remainder compared to the linear rate global convergence guarantee of the Fisher-Rao gradient flow. With general function approximation, it is however not possible to guarantee global convergence [13] and also for other natural gradient methods the linear convergence guarantees are lost when working with function approximation [3, 16] unless one uses regularization. Here, we identify a scenario, where despite being in a function approximation setting, we can ensure global linear convergence.
For a rich enough parametrization Corollary 4.5 ensures that $(\mu_{\theta_{t}})_{t\geq 0}$ follows the Fisher-Rao gradient flow in which case Theorem 3.2 implies the linear convergence of the natural gradient flow (4.1). A common example is the softmax parametrization $\mu_{\theta}(x)\propto e^{\theta(x)}$. For multi-player games with suitable payoff structure, the dynamics of the individual players decouple [14], which allows us to show global convergence for models with exponentially fewer parameters than the softmax parametrization.
@@
Consider a differentiable parametrization of conditional probabilities $\{m_{\theta}:\theta\in\mathbb{R}^{p}\}=\operatorname{int}(\Delta_{\mathbb{X}}^{n})$, where $n\in\mathbb{N}$ and $\mathbb{X}$ is a finite set, and suppose that $\operatorname{span}\left\{\partial_{\theta_{i}}m_{\theta}:i=1,\dots,p\right\}=T\Delta_{\mathbb{X}}^{n}$ for all $\theta\in\mathbb{R}^{p}$. Define a corresponding parametric independence model as
@@ 5. Convergence of State-Action Natural Policy Gradients
They determine the reward as $R(\pi)=\sum_{s\in\mathbb{S},a\in\mathbb{A}}r(s,a)d^{\pi}(s,a)=r^{\top}d^{\pi}$. The set of state-action distributions has been characterized as a polytope, see [20].
@@
where we write $R(\theta)=R(\pi_{\theta})$. Here $G(\theta)$ denotes a Gramian matrix with entries $G(\theta)_{ij}=g_{d_{\theta}}(\partial_{\theta_{i}}d_{\theta},\partial_{\theta_{j}}d_{\theta})$, where we write $d_{\theta}=d^{\pi_{\theta}}$ and $g_{d}$ denotes a Riemannian metric on the state-action polytope $\mathscrsfs{D}$. In this context, the matrix $G(\theta)$ is referred to as a preconditioner. Various choices have been proposed for $G(\theta)$, for example Kakade [26] suggested
@@
which is a weighted sum of Fisher-information matrices over the individual states [26, 9, 46]. This has been studied extensively in the literature, see for example [2, 17, 16, 29], and we refer to it as the Kakade NPG. We focus on the so-called state-action natural policy gradient given by the Fisher information matrix of the state-action distribution [37],
@@
This choice was observed to reduce the severity of plateaus, was used to design a natural actor-critic method [38], and is closely connected to the trust region method known as relative entropy policy search (REPS) [45].
5.1. Convergence guarantees
@@
Consider a finite discounted Markov decision process, suppose 5.2 and 5.3 hold, and consider a solution of the natural policy gradient flow (5.9) for $G=G_{\operatorname{M}}$ and set $R^{\star}\coloneqq\max_{\pi\in\Delta_{\mathbb{A}}^{\mathbb{S}}}R(\pi)$. Then it holds that
@@
  • Expressive exponential families: For a feature map $\phi\colon\mathbb{S}\times\mathbb{A}\to\mathbb{R}^{p}$ and $\theta\in\mathbb{R}^{p}$ we consider the log-linear policy $\pi_{\theta}(a|s)\propto e^{\theta^{\top}\phi(s,a)}$. This provides a regular tabular parametrization if $\operatorname{rank}\{\phi(s,a):s\in\mathbb{S},a\in\mathbb{A}\}=\lvert\mathbb{S}\rvert\cdot\lvert\mathbb{A}\rvert$, see [50, Remark 2.4]. In particular, this includes tabular softmax policies, where $\pi_{\theta}(a|s)\propto e^{\theta_{s,a}}$ which is the arguably most commonly studied policy class.
  • Escort transform: The so-called escort transform $\pi_{\theta}(a|s)\propto\lvert\theta_{s,a}\rvert^{p}$, for a parameter $p\geq 1$ was introduced in [34] to reduce the plateaus of vanilla policy gradients when working with softmax policies.
@@
Consider a finite discounted Markov decision process, suppose 5.2 and 5.3 hold, and consider a solution of the natural gradient flow (5.9) for a regular tabular parametrization and write $\mu_{t}=d_{\theta_{t}}$. Then $(\mu_{t})_{t\geq 0}$ solves the Fisher-Rao gradient flow of the linear program (5.7) and hence Theorem 3.2 and Theorem 3.13 hold. This implies $O(e^{-\Delta t+\kappa\log t})$ convergence for some $\kappa\geq 0$, where
$\begin{split}\Delta=\min\left\{\frac{R^{\star}-R(\pi)}{\lVert d^{\pi^{\star}}-d^{\pi}\rVert_{\operatorname*{\operatorname{TV}}}}:\begin{array}[]{c}\pi\text{ is deterministic and agrees with }\\ \text{a deterministic optimal policy }\pi^{\star}\\ \text{ on all but one state}\end{array}\right\}>0\end{split}$ (5.15)
and $R^{\star}=\max_{\pi\in\Delta_{\mathbb{A}}^{\mathbb{S}}}R(\pi)$ denotes the optimal reward.
@@
The neighbors in $\mathscrsfs{D}$ and $\Delta_{\mathbb{A}}^{\mathbb{S}}$ correspond to each other [41]. Hence, $d^{\pi}$ is a neighbor of $d^{\star}$ if $\pi$ is deterministic and agrees with $\pi^{\star}$ on all but one state. ∎
@@
Much like the state-action natural policy gradient, Kakade’s natural policy gradient with exact gradient evaluations has been shown to converge linearly without the need for entropy regularized setting [29]. Here, the discrete-time setting is studied and NPG is interpreted as soft policy iteration. This is used to show a convergence rate of $R^{\star}-R(\pi_{k})=O(e^{-ck})$ for any $c\in(0,\Delta_{\operatorname{K}})$, where $\Delta_{\operatorname{K}}\coloneqq-(1-\gamma)^{-1}\max\left\{A^{\star}(s,a):a\neq a_{s}^{\star}\right\}\geq\Delta$, where $a^{\star}_{s}$ denotes the optimal action in state $s$. Indeed, by the performance difference lemma, we have
@@
Note that $d\in N(d^{\star})$ can be associated with a policy $\pi$ that agrees with $\pi^{\star}$ on all but one state, and we write $\pi(a_{0}|s_{0})=1$ for $a_{0}\neq a_{s_{0}}^{\star}$ for some $s_{0}$ and $\pi(a_{s}^{\star}|s)=1$ for $s\neq s_{0}$. Since $A^{\star}(s,a_{s}^{\star})=0$ we have $d^{T}A^{\star}=d(s_{0})A^{\star}(s_{0},a_{0})\leq 0$ and estimate
@@
$\Delta\leq-(1-\gamma)^{-1}\max\left\{A^{\star}(s,a):a\neq a_{s}^{\star}\right\}=\Delta_{\operatorname{K}}.$
Hence, the guaranteed convergence rate of Kakade’s NPG is faster compared to the rate we provide for the state-action natural policy gradient. An essentially matching lower bound has been established for Kakade’s NPG in [40], whereas a matching lower bound is missing for state-action natural policy gradients. In our computational example, both converge at the same exponential rate $O(e^{-\Delta_{\operatorname{K}}t})$ even if $\Delta<\Delta_{\operatorname{K}}$.
Remark 5.10 (Implicit bias).
In particular, Corollary 5.8 guarantees that in the case of multiple optimal policies, the gradient flows corresponding to state-action natural policy gradients converge exponentially fast towards the information projection $d^{\star}$ of $d^{\pi_{0}}$ to the set of maximizes $D^{\star}=\{d\in\mathscrsfs{D}:r^{\top}d=R^{\star}\}\subseteq\mathscrsfs{D}$. This shows that state-action natural gradients not only optimize the reward but produce the policy that induces a state-action distribution with maximal entropy with respect to the initial state-action distribution $d^{\pi_{0}}$. This characterizes the implicit bias of state-action natural policy gradients. Prior, the implicit bias of a natural actor-critic method has been analyzed by [25], where they provided an $O(\log k)$ bound on the optimal policy with maximal (weighted) entropy over the states. Note that as this bound grows with the number of iterations $k$ it can’t identify the limiting policy. Further, by using the reformulation as a Hessian gradient flow from [42] and results from convex optimization [5] convergence towards the (generalized) maximal entropy policy for Kakade’s natural policy gradient has been established in [40].
Remark 5.11 (Comparison to previous rates).
In Corollary 5.8 we provide linear convergence with exponent $\Delta$ and have discussed in Remark 3.4 this exponent improves on previously established $O(e-\delta t)$ guarantee in [60, 54]. To see this, we note that $d^{\pi_{1}},d^{\pi_{2}}\in\mathscrsfs{D}$ are neighboring vertices if and only if $\pi_{1},\pi_{2}\in\Delta_{\mathbb{A}}^{\mathbb{S}}$ are neighbors [41]. Two policies are neighboring if and only if they are deterministic and agree on all but one state. Hence, if we consider an MDP with more than one state, there is at least one state $s\in\mathbb{S}$ such that $\pi_{1}(a|s)=\pi_{2}(a|s)=1$ for some $a\in\mathbb{A}$ and therefore $d^{\pi_{1}}(s,a)=d^{\pi_{1}}(s)>0$ and $d^{\pi_{2}}(s,a)=d^{\pi_{2}}(s)>0$. Hence, $d^{\pi_{1}},d^{\pi_{2}}\in\mathscrsfs{D}$ do not have disjoint support and as elaborated in Remark 3.4 this implies $\delta<\Delta$. Overall, this shows that for an exploratory MDP with more than one state, we have $\delta<\Delta$, meaning that our convergence rate improves upon [60, 54].
5.2. Computational examples
We use an example from [26, 9, 37] of an MDP with two states $s_{1},s_{2}$ and two actions $a_{1},a_{2}$, with the transitions and instantaneous rewards shown in Figure 2. We make our code available under https://github.com/muellerjohannes/fisher-rao-GFs-LPs.